Riley company borrowed 36,000 on April 1st year 1 from Titan Bank the note issued by Riley carried a one-year term and a 7% annual interest rate Riley earned cash revenues of 1700 during year one and 1400 during year to assume no other transaction based on this information alone what amount of cash flow from operating activities would appear on the year to statement of cash flows.
Answer #1

Assuming no other transaction during the year,

Cash flow from operating activity:

Net Profit (Loss) during the year(1700-2520) = -820

Add: Interest Paid- to be shown as financing activity = $2,520

Cash flow from Operating Activities = $1,700

Cash flow from Financing activity:

Interest Paid on notes payable = 36,000*7% = ($2,520)

Proceeds from notes issued = $36,000

Similarly in year 2,

Cash flow from OPerating activity would be $1,400